OpenStudy (anonymous):

Suppose that y varies inversely with x, and y = 4 when x = 7. What is an equation for the inverse variation?

OpenStudy (anonymous):

if y varies inversely as x, y=k/x when y=4 and x=7 substituting it into the equation,we have 4=k/7 k=4*7 k=28 the equation for the inverse variation is y=28/x